Saved We have a series RLC circuit with an AC voltage source: The resistance is 100Ohm,...

Saved

We have a series RLC circuit with an AC voltage source: The resistance is 100Ohm, the inductance is 10mH, the capacitance is 10mF. Select all the right answers.

At 60Hz the total impedance is

Question 13 options:

less than 100Ohm

Is more than 100Ohm

Is within 1% of 100 Ohm

Is not within 1% of 100Ohm

Question 14 (1 point)

Saved

We have a series RLC circuit with an AC voltage source: The resistance is 100Ohm, the inductance is 10mH, the capacitance is 10mF. Select all the right answers.

At 60Hz

Question 14 options:

The absolute phase difference between the current and the voltage at the source is less than 4 degree

The absolute phase difference between the current and the voltage at the source is less than 2 degrees

The absolute phase difference between the current and the voltage at the source is less than 6 degrees

The absolute phase difference between the current and the voltage at the source is less than 3 degrees

The absolute phase difference between the current and the voltage at the source is less than 5 degrees

The absolute phase difference between the current and the voltage at the source is less than 1 degree

Question 15 (1 point)

Saved

We have a series RLC circuit with an AC voltage source: The resistance is 100Ohm, the inductance is 10mH, the capacitance is 10mF. Select all the right answers.

At 60Hz what is true?

Question 15 options:

The voltage across the inductor and capacitor lags behind the voltage across the resistor

The voltage across the inductor lags behind the current.

The voltage across the capacitor and resistor lags behind the current

The voltage across the resistor lags behind the voltage across the inductor
Answer #1

13.

Answer is Is B)more than 100Ohm

Inductive reactance

XL=2pifL =2pi*60*10*10-3 = 3.77 ohms

Capacitive reactance

XC =1/2pifC=1/2pi*60*10*10-3 = 0.265 ohms

Impedance

Z=sqrt[R2+(XL-XC)2] =sqrt[1002+(3.77-0.265)2]

Z=100.06 ohms

14)

Answer is D)The absolute phase difference between the current and the voltage at the source is less than 3 degrees

Phase angle

o=tan-1(XL-XC/R) =tan-1[(3.77-0.265)/100]

o=2o

15.

Answer is

The voltage across the inductor and capacitor lags behind the voltage across the resistor